OpenStudy (lena772):

The graph of a system of equations with the same slope and the same y-intercepts will have no solutions. Always Sometimes Never Always right?

ganeshie8 (ganeshie8):

y = mx + c

hero (hero):

@Lena772, you have to test it out

hero (hero):

Trial and error is key here

ganeshie8 (ganeshie8):

yup ! fix m and y intercept and see wat different equations u can get, or if u get any

hero (hero):

@ganeshie8, you should clarify whether or not you're agreeing with me or if you're agreeing with @Lena772

ganeshie8 (ganeshie8):

oh ya, im agreeing wid Hero :)

OpenStudy (anonymous):

If the equations have same slope 'm' then they are parallel. since they all have same y-intercept so the equations should represent a single line more specifically coincident lines...hence the system should have infinite solutions and never 'no solution'

hero (hero):

For example, @Lena772 if you have a system like x + y = 10 5x + 5y = 50 and tried to graph it, what would end up with? No Solution Some Solutions Infinite Solutions

OpenStudy (lena772):

No solution

hero (hero):

Actually...Infinite Solutions

OpenStudy (kc_kennylau):



hero (hero):

The two equations look different, but they are indeed the same. the second one is a multiple of the 1st.

hero (hero):

I simply took x + y = 10 and multiplied both sides by 5 to get 5x + 5y = 50 Multiplying a linear equation in this manner does not change the slope or the y-intercept.

hero (hero):

The way to "see" this is by graphing it. You should take advantage of a site like http://www.desmos.com because it enables you to graph equations quickly. You can definitely use it for graphing systems of equations to compare each equation.